OpenStudy (anonymous):

Write an equation of the line that passes through the point (3,0) and has the slope, m = -2.

OpenStudy (anonymous):

point-slope form is: y-y1 = m(x-x1). m stands for the slope. so you will write y-0 = -2(x-3)

OpenStudy (anonymous):

you just solve that. does that make sense?

OpenStudy (anonymous):

so is it y=mx+3

OpenStudy (anonymous):

yup you just plug-in -2 in place of m...

OpenStudy (anonymous):

oh ok so y=-2x+3

OpenStudy (anonymous):

thks orion :)

OpenStudy (anonymous):

oops i over-looked your intercept...

OpenStudy (anonymous):

it would be a 6 instead of 3

OpenStudy (anonymous):

yup it should be 6...

OpenStudy (anonymous):

oh really y=-2x+6

OpenStudy (anonymous):

thks both of u :)